fairy:

the coffee shop owner explains that you are paid $8/hour and receive a $125 bonus each month for being the manager if the owner hands you a check for $2,501, how many hours did you work this month?write an equation then solve

ThatCreepyBoy:

@Nnesha

ThatCreepyBoy:

@Vocaloid @563blackghost @Shadow

Shadow:

Do you have an idea of how you would approach this?

Shadow:

@fairy

fairy:

nuh

Shadow:

What is our unknown in this problem? What are we trying to solve for?

fairy:

how many hours did i work this month?

Shadow:

Exactly. Let's represent "hours you worked this month" with the variable x Since it tells us "this month" we know that you only worked one month total. Therefore the hours you worked this month, multiplied by 8, plus 125, should come out to 2501.

Shadow:

Can you write an equation for me illustrating that?

fairy:

That's the tricky part ):

Shadow:

Hours you worked "x" times 8 plus 125 equals 2,501

fairy:

8x+125=2,501 ?

Shadow:

Correct, now just solve for the unknown :)

fairy:

how do you do that '-'

Shadow:

Solve for x?

fairy:

ye

Shadow:

You don't know how to solve for x?

fairy:

nuh

Shadow:

Okay, give me a moment to finish something up :)

fairy:

thank you T^T

Shadow:

So have you ever solved for x before...?

Shadow:

Done anything like it?

fairy:

is the answer 297 by any chance?

Shadow:

How did you get that number?

fairy:

2501-125=2377/8=297

Shadow:

Did you do that by yourself?

fairy:

i tried is it wrong? ):

Shadow:

No, it is correct. You know how to solve for x :)

Shadow:

I just wanted to make sure you did it, and know how it works. Solving for x is a foundational concept for Algebra.

fairy:

haha thank you (:

Shadow:

No problem :)